OpenStudy (anonymous):

A point P is placed exactly between two charges, Q1 and Q2. If the electric field experienced by point P due to charge Q1 is 1.5 × 105 newtons/coulomb and the field due to charge Q2 is 7.2 × 105 newtons/coulomb, what is the net electric field at point P?Select one of the options below as your answer: A. 1.5 × 105 newtons/coulomb B. 3.0 × 105 newtons/coulomb C. 5.7 × 105 newtons/coulomb D. 8.7 × 105 newtons/coulomb E. 9.0 × 105 newtons/coulomb

OpenStudy (theeric):

Hi! Have you tried to draw a diagram? A picture's worth 1,000 words or so!

OpenStudy (ivancsc1996):

The net electric field is the sum of both the one caused by Q1 and the one by Q2. If we choose Q2 to be at the right of P and Q1 to be at the left of P, with P being our coordinate system center and the axis x increasing to the right, the equation is E=Q2-Q1=7.2×105-1.5×105=5.7 × 105 newtons/coulomb. So C.